OpenStudy (anonymous):

I have 3 different questions, If you can please help! :) 1. use 9 unit multipliers to convert 10,000 cubic centimeters to cubic miles. 2.A single die is rolled two times. What is the probability that the next roll will produce a number greater then 4? 3. I will Have to draw it out.

OpenStudy (anonymous):

\[\sqrt[4]{24}+ \sqrt{150}+\sqrt[2]{54}\]

OpenStudy (anonymous):